Determine the slope intercept form that passes through (2,3) with equal intercept.
Answer by Alan3354(69443)

Determine the slope intercept form that passes through (2,3) with equal intercept.
------------
If you mean equal x & y intercepts, then the equation is y = -x + b
3 = -2 + b
b = 5
------------
y = -x + 5
x + y = 5